# COLUMBUS: Evaluating COgnitive Lateral Understanding through Multiple-choice reBUSes

### v1

@@ 7 Conclusions
This paper defined a novel visual lateral thinking task, associated with a construction methodology and a resulting synthetic, multiple-choice benchmark COLUMBUS consisting of 1008 rebus puzzles with text and icons. Our experiments on COLUMBUS indicated a substantial gap between human and model performance, which was lessened by adding the graph description to the model inputs. The improvement when expanding a model’s access to include relational rules implies they rely primarily on the elements in a puzzle alone rather than the spatial relationships between them. The analysis showed that models struggle to think laterally about rules that rearrange text, as these require flexible, puzzle-specific abstractions.
The current benchmark’s scale is limited in size. Moreover, its relative coverage of icon and text puzzles and various rules is imbalanced. Future work should apply and extend the methodology provided in this paper to develop more extensive and balanced versions of COLUMBUS, possibly including other multimodal formats beyond rebus puzzles and reducing the variability between puzzle categories to better understand performance gaps. Special attention should be devoted to puzzles that require abstraction and creative thinking and cannot be solved easily by word matching a text/icon. While our methodology includes measures to address this (e.g., homophones and questions with multiple distractors containing the visible words of a puzzle), it can be extended by including synonyms or related words.
@@ A.1 Data Filtering
Of 18,836 puzzles, 4386 remain after automatic filtering (see Appendix A.5). From this subset, we take the top 1000 puzzles according to those that have the most rules per node, plus the most edges (excluding the relational next to rule). We expect these puzzles to be more challenging, as they require one to decode more about an element to understand its use in a puzzle. We manually remove unreadable puzzles that were broken from the generation process, such as too-large elements, overlapping text/icons, and overflowing text/icons. Then, we automatically remove puzzles in which the only rule used is replacing text with icons. As these puzzles are more similar to emoji puzzles than a rebus, we do not consider them to require a sufficient degree of lateral thinking.
@@ A.4 Rule Keywords
A.5 Image Generation
Heuristics.
Our image generation will not render a puzzle in the following four cases:
  1. 1.
    Graphs with nodes connected with the next to relational rule will not be rendered if they have more than three nodes.
  2. 2.
    Graphs that contain an above relational rule will not be rendered if either the top and bottom portions exceed two nodes.
  3. 3.
    Graphs that contain an inside relational rule will not be rendered if either the inside portion has more than one node or the outside portion has more than two nodes.
  4. 4.
    Graphs that contain an outside relational rule will not be rendered if either the inside or outside portions have more than one node.
Templates.
Each template contains $x,y$ coordinates and a multiplier to control the font size. Three of the four templates are points (up to three) spaced roughly equally along the $x$-axis in the middle of the image. Each element in a graph (from left to right) corresponds to a node in the template (the number of nodes in a graph must equal the number of points in a template). The remaining template is used for graphs involving the above relational rule. During image generation, we sequentially fill the points in these templates in with text/icons represented by the graph nodes. We change the appearance of these elements according to the attributes of that element’s respective node.
A.6 Rule Frequency
@@
Modifier Sound 289 150
Repetition: Two 154 77
@@ E.1 Model Results per Prompt
Table 8: Table showing results for four prompts that supply the model with varying levels of information (see Appendix C.1). The results from Table 2 are also shown with prompt 2. Results for Mistral are also reported, but only including the graph description (no image is passed). Highest results in each column are in bold. The models are grouped according to their category outlined in Section 5 (open-source non-instruction VQA, open-source instruction VQA, closed-source VQA, instruction QA).
@@
BLIP-2 OPT (2.7b) 24.8 21.8 18.8 21.5 23.0 24.3 17.0 21.4
BLIP-2 OPT (6.7b) 22.4 24.1 22.9 23.8 26.0 25.7 22.4 21.1
Fuyu (8b) 32.6 32.0 43.0 45.6 32.4 31.1 47.3 49.2
InstructBLIP Vicuna (7b) 50.6 51.3 42.8 41.9 51.6 51.4 38.4 39.2
Qwen-VL (7b) 57.8 57.8 68.0 68.2 68.9 64.3 73.2 72.7
BLIP-2 Flan-T5-XXL (11b) 64.9 68.2 85.1 85.0 67.0 72.2 91.6 90.8
Llava (13b) 63.2 58.0 71.0 71.2 61.4 58.9 69.4 69.2
CogVLM (17b) 59.9 59.3 63.5 64.0 59.5 60.3 64.1 64.9
Llava (34b) 49.2 52.0 62.5 60.3 51.9 49.5 56.3 85.3
GPT-4o 76.4 81.4 90.7 91.0 77.6 84.9 94.4 93.5
GPT-4o-mini 69.0 74.8 88.9 86.7 66.0 77.4 88.0 88.4
Gemini 1.5 (Pro) 65.2 66.5 83.8 84.8 70.5 70.4 89.3 89.8
Gemini 1.5 (Flash) 60.5 61.6 79.9 81.3 63.9 63.4 81.3 81.5
Mistral (7b) N/A N/A 72.3 76.7 N/A N/A 76.8 77.6

### v2

@@ 7 Conclusions
This paper defined a novel visual lateral thinking task associated with a construction methodology and a resulting synthetic, multiple-choice benchmark COLUMBUS consisting of 1005 rebus puzzles with text and icons. Our experiments on COLUMBUS indicated a substantial gap between human and model performance, which was lessened by adding the graph description to the model inputs. The improvement when expanding a model’s access to include relational rules implies they rely primarily on the elements in a puzzle alone rather than the spatial relationships between them. The analysis showed that models struggle to think laterally about rules that rearrange text because these require flexible, puzzle-specific abstractions.
The scale and difficulty of COLUMBUS is still limited. Moreover, its relative coverage of icon and text puzzles and various rules is imbalanced. Future work should apply and extend the methodology provided in this paper to develop more extensive and balanced versions of COLUMBUS, possibly by including other multimodal formats beyond rebus puzzles and reducing the variability between puzzle categories. This can also be improved by increasing the flexibility of the generation process by making graph structures more complex to adjust puzzle difficulty and varying templates across perceptual dimensions like color, positioning, and size. Special attention should also be devoted to puzzles that require abstraction and creative thinking and cannot be solved easily by word matching a text/icon. While our methodology includes measures to address this (e.g., homophones and questions with multiple distractors containing the visible words of a puzzle), it can be extended by including synonyms or related concepts.
@@ A.1 Data Filtering
Of 18,836 puzzles, 4386 remain after automatic filtering (see Section 4). From this subset, we take the top 1000 puzzles according to those that have the most rules per node, plus the most edges (excluding the relational next to rule). We expect these puzzles to be more challenging, as they require one to decode more about an element to understand its use in a puzzle. We manually remove unreadable puzzles that were broken from the generation process, such as too-large elements, overlapping text/icons, and overflowing text/icons. Then, we automatically remove puzzles in which the only rule used is replacing text with icons. As these puzzles are more similar to emoji puzzles than a rebus, we do not consider them to require a sufficient degree of lateral thinking.
@@ A.4 Rule Keywords
A.5 Rule Frequency
@@
Modifier Sound 288 150
Repetition: Two 153 77
@@ E.1 Model Results per Prompt
Table 8: Table showing the mean accuracy (%) averaged across three runs for four prompts that supply the model with varying levels of information (see Appendix C.1). The results from Table 2 are also shown with prompt 2. Results for Mistral are also reported, but only including the graph description (no image is passed). Highest results in each column are in bold. The models are grouped according to their category outlined in Section 5 (open-source non-instruction VQA, open-source instruction VQA, closed-source VQA, instruction QA).
@@
BLIP-2 OPT (2.7b) 24.7 21.9 18.9 21.5 23.2 24.1 16.6 21.1
BLIP-2 OPT (6.7b) 22.2 24.0 23.2 23.6 25.9 25.6 22.6 20.8
Fuyu (8b) 32.4 32.0 42.9 45.5 32.6 31.0 47.4 49.3
InstructBLIP Vicuna (7b) 50.8 51.5 45.3 43.3 51.9 51.8 39.6 38.6
Qwen-VL (7b) 56.7 57.8 67.3 56.0 67.1 63.2 72.0 64.6
BLIP-2 Flan-T5-XXL (11b) 64.8 68.2 85.1 84.9 67.1 72.0 91.6 90.8
Llava (13b) 63.2 58.0 70.9 71.1 61.2 58.8 69.5 69.1
CogVLM (17b) 59.9 59.3 63.5 64.0 59.3 60.1 63.9 64.7
Llava (34b) 67.9 66.8 78.7 79.2 71.3 73.1 85.5 85.4
GPT-4o 75.8 80.9 90.6 90.9 76.4 83.3 93.6 93.2
GPT-4o-mini 67.8 74.0 87.0 86.3 66.4 77.7 88.3 88.9
Gemini 1.5 (Pro) 68.9 71.6 87.5 88.4 72.9 77.5 90.6 91.8
Gemini 1.5 (Flash) 63.6 64.4 86.4 87.0 67.4 67.4 86.7 87.1
Mistral (7b) N/A N/A 70.6 75.3 N/A N/A 73.7 76.8